Lobachevskii Journal of Mathematics

http://ljm.ksu.ru

Vol. 26, 2007, 3–4

IN MEMORY OF BORIS NIKITOVICH SHAPUKOV (15.02.1937 - 13.02.2007)

Boris Nikitovich Shapukov, professor of Kazan State University, was a member of Editorial Board of “Lobachevskii Journal of Mathematics” since the foundation of the journal. He made a distinguished contribution to the journal development actively engaging geometers to submit research papers of highest quality to the journal.

Boris Nikitovich Shapukov was born in Elabuga, Russia. After graduating from secondary school in 1954, he entered mathematical department of Physical and Mathematical Faculty of Kazan State University. Already in his student years B. N. Shapukov showed himself a talented scientific researcher. Under supervision of professor B. L. Laptev, in 1964 B. N. Shapukov defended his thesis ”On geometry of bilinear metric spaces”, where he studied geometry of the affinely connected spaces admitting covariant constant fields of non-symmetric tensors of type (2, 0).

After his defense, developing B L. Laptev’s ideas concerning the theory of ”supporting elements”, B. N. Shapukov starts to study geometry of the general fibered spaces. In his works, B. N. Shapukov develops theory of linear connections and Lie differentiation on the total spaces of smooth fiber bundles. In 1991 B. N. Shapukov defended doctoral thesis ”Structures on fiber bundles and reduction”.

In 1961 B.N. Shapukov was appointed an assistant of the chair of geometry of Kazan State University, in 1976 he became an assistant professor, and in 1992 a professor of the chair of geometry. Generations of students remember his expressive lectures on differential geometry, Lie groups, geometry of fibered spaces, and non-Euclidean geometry.

B. N. Shapukov wrote more than thirty scientific papers, and many tutorials for students, among them ”Tasks on Lie group theory and its applications” (the first edition was published in 1989 by Kazan University, and the second extended edition in 2002 by ”Regular and Chaotic Dynamics”) which was translated into Spanish. He had 7 postgraduates who studied various topics of geometry of fibered spaces.

In 1993 B.N.Shapukov became the head of the chair of geometry. Nineties were rather difficult for Russian science, however at that times B. N. Shapukov organized several international geometrical conferences in Kazan, under his direction ”Proceedings of Geometrical Seminar” were published, he also was the head of a scientific group supported by the Russian government grant program ”Russian Leading Scientific Schools”.

For many years B. N. Shapukov was the Head of International Jury on Lobachevskii Medal awarding, and put many efforts in organizing the work of the jury, as well as international geometrical seminars devoted to this event.

Boris Nikitovich Shapukov will be remembered not only for outstanding contribution to the development of differential geometry, but also for organizing the life of mathematical society.

Editorial Board

Chair of Geometry of Kazan State University